You have a bottle of water. The water has a mass of 35 grams. You heat the water with 125 calories of heat and the final temperature of the water is 42°C. What was the initial temperature (in Celsius) of the water before you added the heat? The specific heat of water is 1 calorie / (°Cg).
Answer #1

use,
q = m*c*(T2-T1)
125 = 35*1*(42 - T1)
(42 - T1) = 3.6
T1 = (42 - 3.6) oC
= 38.4 oC

Answer : 38.4 oC